OpenStudy (anonymous):

guys, can anyone help with this equation..., 24/(2x+6)=3x

OpenStudy (anonymous):

x=1 or x=-4

OpenStudy (anonymous):

here http://www.wolframalpha.com/input/?i=24%2F%282x%2B6%29%3D3x

OpenStudy (anonymous):

24 = 3x(2x+6) 24 = 6x^2 + 18x 6x^2 +18x - 24 = 0 (3x+3)(2x+8) = 0 x = -1, x = -4

OpenStudy (anonymous):

first multiply both sides by (2x+6) to get: 24=6x^2+18x you can use quadratics with c=0

OpenStudy (anonymous):

i mean c=-24

OpenStudy (anonymous):

by using 0=6x^+18x-24

OpenStudy (anonymous):

24/(2x+6)=3x 24=3x(2x+6) 24=6x^2+18x 24=6(x^2+3x) 4 =(x^2+3x) so, x^2+3x-4=0 x^2+4x-x-4=0 (x-1)(x+4)=0 so, x =1 or x =-4

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

:)